OpenStudy (anonymous):

It takes 60oz. of grass seed to seed 3,000 square feet of lawn. At this rate, how much would be needed to seed 5,000 square feet of lawn?

OpenStudy (anonymous):

60s=3000 (x)s=5000 s=(3000/60)=50 50x=5000 x=(5000/50)=100